(x-2)^2=-25 in a+bi form
Answer by stanbon(75887)

(x-2)^2=-25 in a+bi form
----
(x-2)^2+25 = 0
(x-2+5i)(x-2-5i) =0
x = 2-5i or x = 2+5i
========================
Cheers,
Stan H.